QUESTION 5 Australians are one of the highest per capita consumers of coffee in the world. Their coffee consumption has been increasing over the years. As a result, a large number of cafes have continued to open across Brisbane to cater for this increasing demand. Answer the following questions: a. Imagine the market is currently at long run equilibrium. Research has recently been published indicating coffee consumption may lead to increased incidents of cancer. This negative publicity about coffee would be expected to lead to a decline in demand for coffee. As a result, one could expect firms to X for Exit, E for Enter or R for Remain In. the market. Type b. If Marcos café was previously selling 300 coffees per day at a price of $5 per coffee, but is now only selling 230 coffees per day at a price of S4 per coffee. By how much has Marcos economic profit per day declined if his average total costs (which include opportunity cost) were previously $2 per coffee, and are now $3 per coffee? S per day. Answer to the nearest whole number (with no decimal places) c. If Bryans cafe has an average total cost (which includes opportunity cost) of $4.10 per coffee and an average variable cost of $3.90 per coffee, should he keep selling coffees in the short run if the price is $4 per coffee? for No, or M for need More information. Type Y for Yes, N
Answer #1

(a) This negative publicity about coffee would be expected to lead to a decline in demand for coffee. As a result, one could expect firms to exit the market.

(b) When Q=300 and P=$5 , TR=(5)(300)= $1500

When Q=230 and P=$4 , TR = (4)(230)=$ 920

And Earlier AC=$2 , Q=300 , So TC= (2)(300)= $600

So, Earlier Profit = TR-TC = $(1500-600)= $900

After AC =$3 and Q=230 , TC= (3)(230) =$690

So, Profit = $(920-690)= $230

This implies that economic profit declined by $(900-230)= $670

(c) ATC = $4.10

AVC= $3.90

P=$4

Because P>AVC , so Bryan's cafe would not shut down in the short run and continue to produce.

And P<ATC , so it would incur loss in the short run.

YES, he should keep selling coffees in the short run if the price is $4 per coffee.